### v1

# MixFlow: Mixture-Conditioned Flow Matching for Out-of-Distribution Generalization

Abstract
Achieving robust generalization under distribution shift remains a central challenge in conditional generative modeling, as existing conditional flow-based methods often struggle to extrapolate beyond the training conditions. We introduce MixFlow, a conditional flow-matching framework for descriptor-controlled generation that directly targets this limitation by jointly learning a descriptor-conditioned base distribution and a descriptor-conditioned flow field via shortest-path flow matching. By modeling the base distribution as a learnable, descriptor-dependent mixture, MixFlow enables smooth interpolation and extrapolation to unseen conditions, leading to substantially improved out-of-distribution generalization. We provide analytical insights into the behavior of the proposed framework and empirically demonstrate its effectiveness across multiple domains, including prediction of responses to unseen perturbations in single-cell transcriptomic data and high-content microscopy–based drug screening tasks. Across these diverse settings, MixFlow consistently outperforms standard conditional flow-matching baselines. Overall, MixFlow offers a simple yet powerful approach for achieving robust, generalizable, and controllable generative modeling across heterogeneous domains.
33footnotetext: These authors contributed equally to this work.
1 Introduction
Conditional generative models have achieved remarkable success across domains ranging from image synthesis to molecular design (Rombach et al., 2022; Ho and Salimans, 2021). A common paradigm underlying many state-of-the-art approaches, including diffusion models and flow matching, is to learn a transport map from a simple base distribution to a complex target distribution, conditioned on some input descriptor (Lipman et al., 2022; Tong et al., 2023; Benamou and Brenier, 2000). However, a critical limitation emerges when these models encounter conditions not seen during training: their performance degrades substantially, limiting their utility in real-world applications where exhaustive enumeration of all possible conditions is infeasible.
This challenge of out-of-distribution (OOD) generalization***In this manuscript the term OOD generalization refers to the ability of a model to take in an unseen descriptor and accurately generate the corresponding population. This departs from the common notion of the term, that is how a distribution shift in a predictor’s input instances affects its predictions. Moreover, in our setting the unseen descriptor should still come from the same distribution. is particularly acute in scientific domains. Consider drug discovery, where one wishes to predict cellular responses to novel compounds (Subramanian et al., 2017; Srivatsan et al., 2020), or genetic screening, where the goal is to anticipate phenotypic effects of unseen perturbations (Norman et al., 2019; Replogle et al., 2022). In such settings, the space of possible conditions (drugs, genetic perturbations, material compositions) is combinatorially vast, and any practical model must extrapolate beyond its training distribution. Recent work has highlighted that even sophisticated deep learning approaches often fail to outperform simple baselines when evaluated on truly held-out conditions (Ahlmann-Eltze et al., 2025), underscoring the difficulty of this problem.
Why do current conditional generative models struggle with OOD conditions? We argue that the root cause lies in an overlooked architectural choice: the base distribution is fixed. Standard conditional flow matching and diffusion models use a fixed Gaussian $\mathcal{N}(0,I)$ as the source distribution, conditioning only the decoder (velocity field or score network) on the input descriptor (Lipman et al., 2022; Klein et al., 2025; Adduri et al., 2025). This design implicitly treats each condition as an independent estimation problem: the model must learn a separate mapping from the same fixed base to each target distribution. When a novel condition arrives at test time, the model has no mechanism to leverage structural similarities with training conditions; it can only hope that the conditioned decoder generalizes due to, e.g., the decoder being smooth with respect to its input descriptor.
Recent work has begun to explore alternatives to fixed base distributions. Chen et al. (2025) propose modelling the velocity field as a Gaussian mixture to improve few-step sampling, though this does not address the base distribution itself. Atanackovic et al. (2024) introduce Meta Flow Matching, which conditions on initial populations via graph neural network embeddings to model dynamics on the Wasserstein manifold. However, this method works under the assumption that the starting distribution is given to the model and that the nature of the difference between these observed initial distributions is the main source of variability between different effects. In practice, this model is designed to learn how to amortise the effects across contexts rather than generalise to new perturbations. Other approaches design informative priors for specific applications to shorten transport paths and improve sampling efficiency (Bunne et al., 2023). However, none of these methods jointly learn a descriptor-conditioned base distribution with the explicit goal of improving OOD generalization, nor do they provide theoretical justification for why this design choice matters.
Figure 1: Standard conditional flow matching versus MixFlow. (a) Conventional approaches fix the base distribution to a standard Gaussian and condition only the velocity field on the descriptor $y$, effectively learning independent flows for each condition. (b) MixFlow conditions both the base distribution (as a Gaussian mixture) and the velocity field on the descriptor. Similar conditions produce similar base distributions, requiring only small flow corrections and enabling smooth extrapolation to unseen conditions.
We propose a different perspective: the base distribution itself should be conditioned on the descriptor. Our key insight is that when two conditions are similar, whether two rotation angles of images that differ by a small increment, two drugs with related chemical structures (Rogers and Hahn, 2010), or two genetic perturbations affecting the same pathway (Schubert et al., 2018), their corresponding target distributions are also likely to be similar. By learning a base distribution that reflects this similarity structure, we encode prior knowledge about condition relationships directly into the generative process. The flow then needs only to correct for the residual differences, rather than learning the entire mapping from scratch.
We introduce MixFlow, a conditional flow-matching framework that jointly learns a descriptor-conditioned Gaussian mixture base distribution and a descriptor-conditioned velocity field trained via shortest-path (optimal transport) flow matching. The key architectural change is simple: instead of sampling from a fixed Gaussian, MixFlow uses learned networks to predict condition-specific mixture parameters that define the base distribution. Samples from this adaptive base are then transported to the target via a learned velocity field. Because similar descriptors produce similar base distributions, the model can smoothly interpolate and extrapolate to unseen conditions, a capability that fixed-base approaches fundamentally lack.
We provide theoretical analysis supporting this design. Building on connections between flow matching and optimal transport (Peyré et al., 2019; Benamou and Brenier, 2000), we analyze the degrees of freedom in the transport problem as a function of the number of mixture components. We show that when the base is a single Gaussian, the dual optimal transport problem becomes ill-defined, providing a formal explanation for the poor generalization of vanilla conditional flow matching. With multiple mixture components, the problem becomes well-posed, and we derive bounds relating the base complexity, target complexity, and data dimension.
We evaluate MixFlow across diverse domains to demonstrate its broad applicability. On synthetic benchmarks involving rotated letter silhouettes inspired by Atanackovic et al. (2024), MixFlow consistently outperforms conditional flow matching when generalizing to unseen rotations. On image-based phenotypic assays (Caie et al., 2010; Sypetkowski et al., 2023; Bray et al., 2016), MixFlow substantially reduces distributional distances compared to baselines when predicting cellular morphologies under novel compound treatments. On single-cell transcriptomic datasets (Norman et al., 2019; Lotfollahi et al., 2023; Replogle et al., 2022), MixFlow achieves strong performance in predicting gene expression responses to unseen genetic and chemical perturbations.
1.1 Contributions
Our main contributions are:
  • We identify the fixed base distribution as a key limiting factor for OOD generalization in conditional flow matching and propose MixFlow, which jointly learns a descriptor-conditioned Gaussian mixture base and a shortest-path velocity field.
  • We provide theoretical analysis showing that single-Gaussian bases lead to ill-posed inference, while mixture bases with sufficient components yield well-defined problems with bounded estimation error.
  • We demonstrate consistent improvements over conditional flow matching across eight datasets spanning synthetic, image-based, and transcriptomic domains, establishing MixFlow as a flexible framework for conditional generation with improved OOD generalization.
1.2 Related Work
Flow Matching and Continuous Normalizing Flows.
Flow matching has emerged as a scalable framework for training continuous normalizing flows by regressing onto conditional velocity fields along interpolation paths between base and target distributions (Lipman et al., 2022). Unlike score-based diffusion models that require simulating stochastic differential equations, flow matching enables simulation-free training and deterministic sampling via ordinary differential equations. Extensions have incorporated optimal transport couplings to straighten trajectories and improve sample efficiency (Tong et al., 2023; Benamou and Brenier, 2000), while others have explored Riemannian geometries and alternative path constructions (Peyré et al., 2019). A key limitation shared by these approaches is the reliance on a fixed base distribution—typically a standard Gaussian—which we argue fundamentally limits generalization to unseen conditions.
Conditional Generative Models.
Conditioning generative models on auxiliary information has enabled applications from class-conditional image synthesis (Ho and Salimans, 2021; Rombach et al., 2022) to text-to-image generation. In scientific domains, conditional models have been applied to predict cellular responses to perturbations (Lotfollahi et al., 2023; Hetzel et al., 2022; Roohani et al., 2024), model molecular dynamics (Bunne et al., 2023), and generate cell morphologies (Palma et al., 2025; Navidi et al., 2024; Zhang et al., 2025b). Recent flow-based approaches include CellFlow (Klein et al., 2025), which applies flow matching to single-cell phenotype modeling, and STATE (Adduri et al., 2025), a transformer-based architecture for context-aware perturbation prediction. These methods condition the velocity field or decoder on perturbation descriptors but maintain a fixed base distribution, treating each condition as an effectively independent estimation problem. Despite their success in different domains, generative models generalization remains a puzzling phenomenon Bertrand et al. (2025). For example, a line of research shows that generative models may learn to merely generative training instances, failing to generate novel samples and risking privacy issues when trained on licenced or sensitive medical data Yoon et al. (2023); Bertrand et al. (2025); Bamberger et al. (2025). Besides the memorisation viewpoint, one can see the lack of generalisation from the lens of score estimation Song and Ermon (2019), where the generative model is shown to learn an independent score estimator for each descriptor in the training set and its corresponding population, as well as for the combined population Ho and Salimans (2021). In this case, the independent score estimators are not expected to generalise to unseen descriptors.
Gaussian Mixtures in Generative Models.
Gaussian mixture models (GMMs) have a long history as flexible density estimators with universal approximation properties (Nguyen, 2019; Zheng and Tang, 2020). Recent work has revisited GMMs in the context of modern generative models. Chen et al. (2025) propose GMFlow, which models the velocity distribution as a Gaussian mixture to capture multimodal flow directions, improving few-step sampling and reducing over-saturation artifacts. Ahamed et al. (2025) use GMMs in latent spaces for molecular generation. These approaches modify how the velocity or score is represented but do not condition the base distribution on input descriptors. In contrast, MixFlow learns a descriptor-conditioned GMM as the source distribution, explicitly encoding condition similarity structure to improve OOD generalization.
Learned and Adaptive Prior Distributions.
Several works have explored moving beyond fixed Gaussian priors. In the context of variational autoencoders, learned priors have been used to improve expressiveness (Bereket and Karaletsos, 2023). For flow-based models, Bunne et al. (2023) learn transport maps via neural optimal transport for single-cell perturbation modeling, though with a focus on paired control-treatment settings rather than OOD generalization. Meta Flow Matching (Atanackovic et al., 2024) conditions on initial populations by embedding them via graph neural networks, enabling generalization across different starting distributions on the Wasserstein manifold. This approach is complementary to ours: while Meta Flow Matching learns to adapt to varying initial populations, MixFlow learns to predict appropriate base distributions from condition descriptors, directly targeting the OOD generalization problem.
Out-of-Distribution Generalization.
OOD generalization remains a central challenge in machine learning, with extensive work on domain adaptation, invariant learning, and distributionally robust optimization. In the context of perturbation biology, recent analyses have shown that deep learning models often fail to generalize beyond training perturbations, sometimes underperforming simple linear baselines (Ahlmann-Eltze et al., 2025). This has motivated the development of methods that explicitly encode biological structure, such as pathway information (Roohani et al., 2024) or chemical similarity (Subramanian et al., 2017; Wang et al., 2016). MixFlow takes a complementary approach by encoding similarity structure directly into the generative model architecture through the learned base distribution, providing a general mechanism for OOD generalization that does not require domain-specific inductive biases.
2 Problem Setup
We study population-level generative modelling. Let $\{\rho_{1},\ldots,\rho_{N}\}$ denote $N$ probability measures on $\mathbb{R}^{D}$, each corresponding to a distinct population. For each $n\in\{1,\ldots,N\}$ we observe a dataset
$\bm{X}_{n}=\{\bm{x}_{n,1},\ldots,\bm{x}_{n,s_{n}}\}\subset\mathbb{R}^{D},$
consisting of $s_{n}$ samples drawn i.i.d. from $\rho_{n}$. Each population is associated with a descriptor $\bm{y}_{n}\in\mathcal{Y}$ that encodes metadata about the population (e.g., the rotation angle of a letter, perturbation applied to cells, a molecular embedding of a drug, or cell-type information).
Given training data $\{(\bm{X}_{n},\bm{y}_{n})\}_{n=1}^{N}$, the aim is to learn a conditional generative model that produces samples from $\rho_{n}$ when provided with $\bm{y}_{n}$. At test time, the model is presented with an unseen descriptor $\bm{y}_{\mathrm{test}}$ and must generate samples approximating the corresponding, unknown distribution $\rho_{\mathrm{test}}$. This setup captures, for example, predicting the distribution of gene expression states under new perturbations, conditioned on information about the perturbation, baseline cells, and biological experimental conditions.
2.1 Background on Conditional Flow Matching
Flow matching provides a scalable framework for learning conditional generative models that map a simple base distribution to a potentially-complex one.
Flow-based generative models.
A flow model generates a target distribution $\rho$ by transporting a simple base distribution $\mu$ (e.g. a standard Gaussian) along the trajectories of an ODE. Let $v:\mathbb{R}^{D}\times[0,1]\to\mathbb{R}^{D}$ be a time-dependent velocity field and consider
$\frac{d}{dt}\bm{x}_{t}\;=\;v(\bm{x}_{t},t),\qquad\bm{x}_{0}\sim\mu.$ (1)
The measure on $\bm{x}_{t}$ is denoted by $\rho_{t}$, which evolves according to the continuity equation:
$\partial_{t}\rho_{t}\;=\;-\nabla\cdot(\rho_{t}v(\cdot,t)),\qquad\rho_{0}=\mu.$ (2)
With an appropriate choice of $v$, the terminal distribution $\rho_{1}$ coincides with the target $\rho$.
Conditional flows.
To capture population variability, we parameterize a conditional velocity field
$v_{\theta}:\mathbb{R}^{D}\times[0,1]\times\mathcal{Y}\to\mathbb{R}^{D},$
and define the dynamics
$\frac{d}{dt}\bm{x}_{t}\;=\;v_{\theta}(\bm{x}_{t},t;\bm{y}),\qquad\bm{x}_{0}\sim q.$ (3)
For a descriptor $\bm{y}_{n}$, the terminal distribution of this flow, denoted $\rho_{1,\theta}(\cdot\mid\bm{y}_{n})$, should approximate $\rho_{n}$. At test time, for a novel descriptor $\bm{y}_{\mathrm{test}}$, the model produces $\rho_{1,\theta}(\cdot\mid\bm{y}_{\mathrm{test}})$ as an approximation to $\rho_{\mathrm{test}}$.
Flow matching objective.
Training $v_{\theta}$ is performed by regression to a teacher velocity field that imposes the boundary conditions (i.e. the conditions $\rho_{0}=\mu$ and $\rho_{1}=\rho$) along with potentially other conditions to, e.g., regularise the flow model. The teacher flow is constructed by integrating over interpolant paths between the base $\mu$ and the empirical populations $\rho_{n}$, while the training is shown to be feasible by individual interpolant paths Lipman et al. (2022). Specifically, let $(\bm{x}_{0},\bm{x}_{1})\sim\mu\otimes\rho_{n}$, i.e. $\bm{x}_{0}\sim\mu$ and $\bm{x}_{1}\sim\rho_{n}$ independently, and consider the linear interpolant
$\bm{x}_{t}\;=\;(1-t)\bm{x}_{0}+t\bm{x}_{1},\qquad t\in[0,1].$
The individual conditional teacher interpolant path is defined as
$u(\bm{x}_{t},t\;|\bm{x}_{0},\bm{x}_{1})\;=\;\bm{x}_{1}-\bm{x}_{0}.$ (4)
More advanced constructions use Wasserstein geodesic interpolants (Benamou and Brenier, 2000; Peyré et al., 2019; Tong et al., 2023), which provide velocity fields aligned with optimal transport.
The conditional flow matching loss is then
$\mathcal{L}(\theta)\;=\;\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{t\sim\mathrm{Unif}[0,1]}\mathbb{E}_{(\bm{x}_{0},\bm{x}_{1})\sim\mu\otimes\rho_{n}}\left\|v_{\theta}\!\big((1-t)\bm{x}_{0}+t\bm{x}_{1},t;\bm{y}_{n}\big)-\big(\bm{x}_{1}-\bm{x}_{0}\big)\right\|^{2}.$ (5)
Minimizing equation 5 learns a parametric family of velocity fields that transports the base distribution $\mu$ to each $\rho_{n}$, while sharing parameters across populations through conditioning on $\bm{y}_{n}$.
3 MixFlow: Mixture-Conditioned Flow Matching
Standard conditional flow matching fixes the base distribution to $\mathcal{N}(\mathbf{0}_{D},\sigma^{2}\mathbf{I}_{D\times D})$ and conditions only the decoder, effectively yielding independent estimators for each population and limiting generalization (See, Fig. 1(a)). MixFlow instead conditions both the base distribution and the flow module on perturbations and covariates (Fig. 1(b)).
Let $I$ denote the number of mixture components. Plausible base distributions are Gaussian mixtures $\mu_{\mathrm{GMM}(\bm{\Theta},\bm{p})}$ with $\bm{\Theta}\in\mathbb{R}^{I\times D}$ specifying mode locations, $\bm{p}\in\mathbb{S}^{I-1}$ specifying mixture weights, and fixed component variance $\sigma^{2}$. The family of such mixtures forms the restricted subspace illustrated in Fig. 1(b) in orange.
Given populations $\{\rho_{1},\ldots,\rho_{N}\}$ with descriptors $\{\bm{y}_{1},\ldots,\bm{y}_{N}\}$, each $\rho_{n}$ is projected to its closest element $\mu_{\mathrm{GMM}(\bm{\Theta}^{(n)},\bm{p}^{(n)})}$ in this subspace. The optimal transport to $\rho_{n}$ from its projection defines a velocity field $v(.,t;\bm{y}_{n})$. These velocity fields form geodesics on the 2-Wasserstein manifold of probability measures, which are shown by the red curves in Fig. 1(b). Predictors $h_{\bm{\Theta}}:\mathcal{Y}\to\mathbb{R}^{I\times D}$ and $h_{\bm{p}}:\mathcal{Y}\to\mathbb{S}^{I-1}$ are trained so that $h_{\bm{\Theta}}(\bm{y}_{n})\approx\bm{\Theta}^{(n)}$ and $h_{\bm{p}}(\bm{y}_{n})\approx\bm{p}^{(n)}$. At test time, the predicted base $\mu_{\mathrm{GMM}(h_{\bm{\Theta}}(\bm{y}_{test}),h_{\bm{p}}(\bm{y}_{test}))}$ is evolved by $v(.,t;\bm{y}_{test})$ to approximate $\rho_{test}$.
3.1 Training objectives
MixFlow is trained with two complementary losses.
Objective 1. Align the learned velocity field $v_{\theta}(.,t;\bm{y}_{n})$ with the OT path between $\rho_{n}$ and its projected base:
$\mathcal{L}_{\mathrm{OT}}=\mathbb{E}\Bigg[\sum_{s=1}^{S}\big\|v\!\left(t,(1-t)\bm{x}^{(0)}_{s}+t\bm{x}_{\tau(s)},\bm{y}_{n}\right)-(\bm{x}_{\tau(s)}-\bm{x}^{(0)}_{s})\big\|^{2}\Bigg],$ (6)
where $\{\bm{x}_{s}\}\sim\rho_{n}$, $\{\bm{x}^{(0)}_{s}\}\sim\mu_{GMM(h_{\bm{\Theta}}(\bm{y}_{n}),h_{\bm{p}}(\bm{y}_{n}))}$, $t\sim\mathrm{Unif}(0,1)$, and $\tau$ denotes mini-batch OT pairing (Tong et al., 2023).
Objective 2. Encourage each projection $\mu_{GMM(h_{\bm{\Theta}}(\bm{y}_{n}),h_{\bm{p}}(\bm{y}_{n}))}$ to remain close to $\rho_{n}$ by minimizing geodesic length:
$\mathcal{L}_{\mathrm{geo}}=\mathbb{E}\Bigg[\sum_{s=1}^{S}\big\|\bm{x}_{\tau(s)}-\bm{x}^{(0)}_{s}\big\|^{2}\Bigg],$ (7)
with the Gumbel–softmax trick ensuring differentiability with respect to mixture weights and mode locations.
The training procedure is elaborated in Alg. 1, where the training alternates between updating the flow model in line 19 and the base distributions in line 22. In Alg. 1 the training planner $P$ divides the training into three parts: (i) warm-up period, where only the flow model is updated, (ii) alternating period, during which after a certain number of updates to the flow model, a single update is made to the base distributions. (iii) cool-down period, where the base distributions are kept fixed and only the flow model is trained. Notably, during the warm-up and alternating periods the dropout layers of $H_{\bm{\Theta}}$ and $H_{\bm{p}}$ are enabled, while in the cool-down period they are disabled.
3.2 Inference
At test time, given a novel descriptor $\bm{y}_{test}$, MixFlow predicts the base distribution $\mu_{\mathrm{GMM}(h_{\bm{\Theta}}(\bm{y}_{test}),h_{\bm{p}}(\bm{y}_{test}))}$ and evolves it with $v(.,t;\bm{y}_{test}),\;t\in[0,1]$ to generate samples approximating $\rho_{test}$.
Algorithm 1 Training MixFlow with mixture-conditioned flow matching
1:Datasets $\{(\bm{X}_{n},\bm{y}_{n})\}_{n=1}^{N}$, number of base GMM modes II, fixed GMM variances $\sigma^{2}$, training planner $P$.
2:Initialize parameters: $params(v(.,.,.))$ (velocity field), $params(h_{\bm{\Theta}}(.))$ (mode position predictor), and $params(h_{\bm{p}}(.))$ (mode probability predictor), learning rate $\ell$.
[…]
5:for each training epoch do
6:  while the epoch hasn’t finished do
7:   $mode\_train,\;flag\_settoeval\_H\leftarrow P.next(cnt\_epoch,\;cnt\_iteration)$
8:   Sample a population $(\bm{X}_{n},\bm{y}_{n})$
9:   $\{\bm{x}_{s}\}_{s=1}^{S}\leftarrow$ $S$ samples from the selected population
10:   $\{\bm{x}^{(0)}_{s}\}_{s=1}^{S}\leftarrow S\;\text{Samples from }\mu_{GMM(h_{\bm{\Theta}}(\bm{y}_{n}),h_{\bm{p}}(\bm{y}_{n}))}$
11:   $\tau\leftarrow$ mini-batch discrete OT pairing between $\{\bm{x}_{s}\}_{s=1}^{S}$ and $\{\bm{x}^{(0)}_{s}\}_{s=1}^{S}$
12:   if flag_settoeval_Hflag\_settoeval\_H then
13:     $H_{\bm{\Theta}}.eval()$;  $H_{\bm{p}}.eval()$ $\triangleright$ i.e. disable dropout
14:   else
15:     $H_{\bm{\Theta}}.train()$;  $H_{\bm{p}}.train()$ $\triangleright$ i.e. enable dropout
16:   end if
17:   if $mode\_train\text{\; equals 'velocity field'}$ then
18:     $t\leftarrow unif(0,1)$
[…]
20:     $params(v(.,.,.))\leftarrow params(v(.,.,.))-\ell\cdot\nabla\mathcal{L}_{OT}$
21:   else if $mode\_train\text{\; equals 'base distribution'}$ then
[…]
23:     $params(H_{\bm{\Theta}},H_{\bm{p}})\leftarrow params(H_{\bm{\Theta}},H_{\bm{p}})-\ell\cdot\nabla\mathcal{L}_{geo}$
24:   else
25:     "print(ERROR: unknown value for mode_trainmode\_train)"
26:   end if
27:   $cnt\_iteration\leftarrow cnt\_iteration+1$
28:  end while
29:  $cnt\_epoch\leftarrow cnt\_epoch+1$
30:end for
4 Generalization Analysis
The Wasserstein manifold of probability measures is infinite-dimensional, hindering the adoption of theoretical generalization guarantees for finite-dimensional spaces. To circumvent this, we assume that the target distribution $\rho$ is itself a Gaussian mixture model (GMM) with $J$ components,
$\rho=\rho_{GMM(\bm{\Gamma},\bm{q})},$
where $\bm{\Gamma}\in\mathbb{R}^{J\times D}$ are fixed mode locations, $\bm{q}\in\mathbb{S}^{J-1}$ are mixture weights, and all components share a fixed variance. This setting still captures highly general distributions: as $J\to\infty$ and $\bm{\Gamma}$ forms a sufficiently fine grid covering $\mathbb{R}^{D}$, the family $GMM(\bm{\Gamma},\bm{q})$ is dense in the space of probability measures under the Wasserstein metric (Nguyen, 2019; Zheng and Tang, 2020). Thus, assuming $\rho$ is GMM-induced is a broad and flexible modelling assumption. In this formulation, prediction reduces to estimating the mixture weights $\bm{q}\in\mathbb{S}^{J-1}$ from the population descriptor $\bm{y}$.
A naive solution is to directly learn a predictor mapping $\bm{y}\mapsto\bm{q}$. However, this approach suffers from poor generalization as $J$ grows, since the dimensionality of $\bm{q}$ directly affects the generalization. MixFlow instead takes a structured approach: rather than predicting $\bm{q}$ directly, it (i) projects the target distribution onto the subspace of GMMs with $I$ components (possibly, $I\ll J$), and (ii) learns the velocity field that transports this projection to the target distribution.
Besides the aforementioned assumptions, in the theoretical analysis we assume that this 2-step algorithm uses mixture Wasserstein distance (Delon and Desolneux, 2020; Bing et al., 2022) and the flow model induced by it, as opposed to the commonly used 2-Wasserstein distance. In the latter case, the pushforward function of the flow model can map its input points to any arbitrary point in $\mathbb{R}^{D}$. While in the former case, a discrete optimal transport problem is solved between the source and target GMMs to determine, intuitively, how the modes of the source GMM should be moved or split to create the target GMM, and that fully determines the flow. In this sense, the latter metric and flow model are a generalisation of the former. Nonetheless, this analysis highlights the importance of conditioning the base distribution on population descriptor $\bm{y}$, which we experimentally validate in in Sec. 6.
In this section we only summarize the key theoretical findings. Details are provided in Appendix A.
Statement 1 (Summary of theoretical findings).
Recall that $I,J,D\in\mathbb{N}$ are respectively the number of modes in the base and target GMM distributions, and the data dimension.
  1. §1
    In the testing phase, the velocity field $\bm{V}\in\mathbb{R}^{I\times J}$ is identified up to JIDJ-ID degrees of freedom. In particular, if $I\geq\lceil J/D\rceil$, then $\bm{V}$ is uniquely identified.
  2. §2
    If $I=1$, the problem reduces to predicting the $J$ mixture weights $\bm{q}$ directly from $\bm{y}$, corresponding to the common baseline of fixing the base distribution to $\mathcal{N}(\bm{0}_{D},\bm{I}_{D\times D})$ and explaining its poor generalization.
Essentially, Statement 1 suggests that increasing the number $I$ of modes of the base distribution by one, reduces degrees of freedom of the flow model by the data dimension $D$, thereby facilitating the prediction and identification of the flow model in the testing phase.
5 Experiments
We evaluate MixFlow against Vanilla Conditional Flow Matching (Vanilla CFM) on a set of synthetic and experimental datasets. Our goal is to assess MixFlow’s generalizability and its robustness across multiple conditions, datasets, and hyperparameters. We seek empirical evidence of the need for the Gaussian Mixture Model (GMM) to describe the base distribution and its impact on the distributional distance between the inferred final distributions and the empirical ones.
5.1 MixFlow on Synthetic Data
Following Meta Flow Matching (Atanackovic et al., 2024), we create a synthetic benchmark of populations $\{(\rho_{i},\bm{y}_{i})\}$ where each condition $i$ corresponds to a target population $\rho_{i}=p_{1}(x_{1}\mid i)$ supported on a rendered letter silhouette. For every alphabet character we generate 20 rotations; each (letter, rotation) pair defines one condition (i.e. one descriptor $\bm{y}_{i}$) and its corresponding population. The descriptor $\bm{y}_{i}$ is the concatenation of a one-hot letter code and a normalized rotation value. The target population $\rho_{i}$ is formed by sampling uniformly from the foreground pixels of the silhouette. To ensure that each condition corresponds to a full population rather than a single sample, we render multiple copies of the same silhouette using different fixed RGB colours. These colour variations are used only to generate several samples per condition and are not included in the descriptor $\bm{y}_{i}$.
For MixFlow, we construct a mixture-conditioned base for each population. We set the number of base modes $I$ equal to the number of conditions and assign one Gaussian component per condition. In our synthetic setup, we do not use $h_{\bm{\Theta}}(\bm{y}_{i})$ to predict the component means; instead, the GMM mode locations are implemented as free learnable parameters, each initialized from a single rendered sample of its corresponding condition. The mixture weights are still predicted by the MLP $h_{\bm{p}}(\bm{y}_{i})$. The velocity field $v(\cdot,t;\bm{y}_{i})$ is then trained using the flow-matching objectives in Eqs. 67 to transport the resulting $\mu_{\mathrm{GMM}(\bm{\Theta}^{(i)},h_{\bm{p}}(\bm{y}_{i}))}$ toward $\rho_{i}$.
As a baseline, CFM fixes the base to $\mathcal{N}(\mathbf{0}_{D},\mathbf{I}_{D\times D})$ and samples $x_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at generation time. We train both MixFlow and CFM on a synthetic dataset of 6 letters with $R=20$ rotations per letter, and study rotation interpolation generalization. Specifically, we generate $R$ rotation angles uniformly spaced in $0,2\pi]$ and split them by index: the training set contains the 10 even-indexed rotations, while the held-out set contains the 10 odd-indexed rotations. The validation set is formed by selecting a single val_letter and evaluating that letter only on the held-out (odd-indexed) rotations; all letters (including val_letter) are present in training but only at the even-indexed rotations. We report two independent runs with identical settings except for the choice of validation letter, using $\texttt{val\_letter}\in\{\texttt{S},\texttt{H}\}$. We do this to check that conclusions are not tied to a single character’s geometry and to reduce sensitivity to a particular letter choice. In each run, the held-out set contains 10 unseen rotations (odd indices) for the chosen validation letter (S or H). Results are reported in Table 1. Qualitative samples are available in Figure 2.
We evaluate both models using three standard distributional metrics: Wasserstein-1 (W1), Wasserstein-2 (W2), and Maximum Mean Discrepancy (MMD). For each experiment, we select the best checkpoint based on validation Wasserstein-2 distance and report all metrics at this checkpoint. Across both validation-letter settings (S and H), MixFlow consistently improves over conditional flow matching (CFM) in all three metrics. In particular, MixFlow achieves lower W1 (reducing by $\sim 39\%$ and $\sim 13\%$), lower W2 (reducing by $\sim 16\%$ and $\sim 13\%$), and lower MMD (reducing by $\sim 17\%$ and $\sim 12\%$) for both experiments, demonstrating that mixture-conditioned base distributions provide a more expressive and accurate starting point than the fixed isotropic Gaussian used by CFM.
Figure 2: Qualitative comparison of Conditional Flow Matching (CFM) and MixFlow on the synthetic letter–rotation benchmark. Each block shows samples at different integration times ($t=0$, $t=0.5$, $t=1$) and the corresponding target distribution. MixFlow produces smoother trajectories and more faithful reconstructions on both training and unseen rotations, while CFM exhibits drift and shape distortion, particularly on validation conditions.
Table 1: Results on two synthetic experiments based on the letter S and letter H silhouettes. Each experiment assesses a model’s ability to generalize to rotations of the corresponding shape that are not observed during training.
Dataset Model MMD $\downarrow$ W1 $\downarrow$ W2 $\downarrow$
Letter S CFM 0.689 $\pm$ 0.0140 0.01533 $\pm$ 0.00094 0.5667 $\pm$ 0.0114
MixFlow 0.5737 $\pm$ 0.0071 0.00933 $\pm$ 0.00047 0.4723 $\pm$ 0.0057
Letter H CFM 0.8363 $\pm$ 0.0193 0.1857 $\pm$ 0.0061 0.6777 $\pm$ 0.0168
MixFlow 0.7297 $\pm$ 0.0061 0.1380 $\pm$ 0.0016 0.5867 $\pm$ 0.0024
5.2 Using MixFlow to predict single-cell responses for unseen interventions
Single-cell perturbation screenings are an invaluable tool for new treatment discovery; they consist in testing the effects of genetic or chemical perturbations on cells in vitro. Each cell is exposed to a specific condition, and its gene expression is measured to assess whether the perturbation is associated with a given phenotypic change. We evaluate whether MixFlow can predict cellular responses to perturbations never observed before.
Models are trained on a subset of conditions and tested on held-out ones, defining an out-of-distribution (OOD) generalisation problem. Table 2 provides an overview of the datasets used in our experiments. Notably, Combo-Sciplex (Lotfollahi et al., 2023) and Norman (Norman et al., 2019) datasets also contain combinations of two perturbations. For each dataset, we generate four random splits, holding out 70% of the perturbations from the training set. Given the larger number of cells available in the iAstrocytes and Replogle-Nadig (Replogle et al., 2022) datasets, we guarantee high reliability of our results by filtering for conditions with at least 300 and 500 cells, respectively. The iAstrocytes data has been produced in-house in Sanger Institute, and it is not currently publicly available.
We report Wasserstein-1, Wasserstein-2, MMD, and Energy distances (ED). These metrics assess whether predicted distributions accurately capture the global geometry of observed cell states, beyond low-order statistics. On all tested single-cell perturbation datasets, MixFlow achieves substantially stronger distributional alignment than Vanilla CFM (Table 3), showing how MixFlow better generalises when generating previously unseen conditions, across both genetic and chemical perturbation modalities.
Table 2: Single-cell perturbation datasets used to evaluate MixFlow.
Dataset Cells Perturbations Cell type
Norman (Norman et al., 2019) $\sim$90k 277 CRISPR K562
Combo-Sciplex (Lotfollahi et al., 2023) $\sim$70k 32 compounds A549
iAstrocytes $\sim$150k 148 CRISPR iAstrocytes
Replogle–Nadig (Replogle et al., 2022) $\sim$250k 67 CRISPRi Jurkat
BBBC021 (Caie et al., 2010) 97,504 images 35 compounds MCF7
RxRx1 (Sypetkowski et al., 2023) 170,193 images 1066 siRNA U2OS
single or paired perturbations.
Table 3: Cell-type perturbation benchmarks comparing MixFlow and CFM. Mean $\pm$ std over four runs. Lower is better.
Dataset Model MMD $\downarrow$ W1 $\downarrow$ W2 $\downarrow$ ED $\downarrow$
Norman CFM 0.199 $\pm$ 0.009 30.449 $\pm$ 0.241 467.8 $\pm$ 8.1 16.2 $\pm$ 0.3
MixFlow 0.192 $\pm$ 0.004 30.084 $\pm$ 0.284 456.3 $\pm$ 9.1 15.9 $\pm$ 0.3
Combo-SciPlex CFM 0.141 $\pm$ 0.005 26.007 $\pm$ 0.464 344.8 $\pm$ 12.6 9.0 $\pm$ 0.3
MixFlow 0.129 $\pm$ 0.006 25.640 $\pm$ 0.180 335.3 $\pm$ 4.8 8.5 $\pm$ 0.2
iAstrocytes CFM 0.163 $\pm$ 0.003 27.121 $\pm$ 0.110 369.7 $\pm$ 3.0 13.2 $\pm$ 0.1
MixFlow 0.135 $\pm$ 0.004 27.061 $\pm$ 0.233 367.6 $\pm$ 6.5 12.6 $\pm$ 0.1
Replogle (Jurkat) CFM 0.116 $\pm$ 0.006 52.919 $\pm$ 0.200 1408.3 $\pm$ 10.6 28.0 $\pm$ 0.3
MixFlow 0.095 $\pm$ 0.003 52.868 $\pm$ 0.215 1404.9 $\pm$ 11.5 27.3 $\pm$ 0.4
5.3 MixFlow predicts unseen intervention responses in image-based assays
To extend beyond transcriptomic data, we evaluate MixFlow in an image-based drug discovery setting. High-content imaging provides rich phenotypic readouts, but exhaustive compound testing is infeasible. We therefore train on a subset of compounds and predict cell morphologies for unseen compounds using only their chemical representations at test time.
We benchmark on BBBC021 Caie et al. (2010) and RxRx1 Sypetkowski et al. (2023), comparing MixFlow against a conditional flow matching baseline (CFM). Table 4 shows that MixFlow achieves consistently lower distance metrics (MMD, W1, W2, ED), indicating more faithful recovery of unseen phenotypic responses. We split BBBC021 following prior work (Palma et al., 2025): training uses 25 compounds and testing 9. For RxRx1, we train on 853 siRNAs and test on 213.
A key design choice is to generate instances in feature space rather than pixel space. Images are encoded with DINOv2 (Oquab et al., 2024) into $768$-D embeddings, which aligns with semantically meaningful morphology-level distances (Oquab et al., 2024; Zhang et al., 2018). This strategy improves efficiency, imposes a semantic inductive bias, and enables geometry-aware evaluation. For fairness, CFM outputs are also scored in the same DINOv2 feature space. Results are provided in Table 4. These substantial gains indicate that MixFlow more faithfully recovers unseen compound phenotypes, highlighting its potential to reduce experimental burden through reliable in silico prediction.
Table 4: BBBC021 and RxRx1 benchmarks comparing MixFlow and CFM on phenotypic perturbations. Lower is better; values are the mean over three runs.
Dataset Model MMD $\downarrow$ W1 $\downarrow$ W2 $\downarrow$ ED $\downarrow$
BBBC021 CFM 0.2504 $\pm$ 0.001 108.4093 $\pm$ 0.291 5903.8647 $\pm$ 32.33 18.9197 $\pm$ 0.088
MixFlow 0.0439 $\pm$ 0.002 29.4752 $\pm$ 0.733 442.9421 $\pm$ 24.11 2.4036 $\pm$ 0.087
RxRx1 CFM 0.2527 $\pm$ 0.001 118.7191 $\pm$ 1.23 7077.8330 $\pm$ 43.21 20.5810 $\pm$ 0.31
MixFlow 0.0307 $\pm$ 0.001 30.8660 $\pm$ 0.12 483.2796 $\pm$ 4.69 1.6731 $\pm$ 0.034
6 Ablations
We conduct ablation studies to understand the key design choices in MixFlow. All experiments are performed on both transcriptomic (iAstrocytes) and morphological (BBBC021) datasets to ensure findings generalize across modalities.
Number of mixture components.
A central design choice in MixFlow is the number of modes $I$ in the Gaussian mixture base distribution. We vary $I$ (the number of modes) and measure the energy distance between generated and ground-truth distributions on held-out perturbations. As shown in Figure 3(a), increasing the number of modes consistently improves performance up to a point: energy distance decreases as $I$ grows, reflecting the increased expressiveness of the base distribution. However, beyond a dataset-dependent threshold, performance plateaus or slightly degrades. We attribute this to optimization difficulty: with too many modes, the model must learn to allocate probability mass across components that may not be necessary, increasing the effective number of parameters without proportional benefit. Notably, the optimal $I$ differs between datasets (transcriptomic data benefits from more modes than morphological data) suggesting that the appropriate complexity of the base distribution should reflect the heterogeneity of the target populations. These findings align with our theoretical analysis in Section 4: increasing $I$ reduces the degrees of freedom in the transport problem by $D$ per mode, improving identifiability (Statement 1, §2).
Robustness to data dimensionality.
A known challenge for flow-based models is scaling to high-dimensional spaces, where the fixed Gaussian base distribution $\mathcal{N}(0,I)$ becomes increasingly mismatched to structured target distributions. We investigate this by training both MixFlow and CFM on PCA-reduced representations with increasing numbers of components. Figure 3(b) shows that CFM’s performance degrades substantially as dimensionality increases: the energy distance grows by over an order of magnitude between 50 and 500 PCA components. In contrast, MixFlow maintains stable performance across the full range of dimensionalities tested. This robustness arises because the learned GMM base distribution adapts to the structure of each condition, reducing the effective transport distance even in high dimensions. From the perspective of Section 4, higher $D$ increases the constraint count ($ID+I$), which for fixed $I$ reduces the residual degrees of freedom JIDJ-ID, explaining why MixFlow’s advantage becomes more pronounced in higher dimensions.
These ablations confirm two key properties of MixFlow: (i) the number of mixture components provides a useful complexity knob that should be tuned to dataset heterogeneity, and (ii) conditioning the base distribution yields substantial robustness benefits in high-dimensional settings where standard CFM struggles.
(a) Energy distance versus number of GMM modes on transcriptomic (left) and morphological (right) datasets. Performance improves with more modes up to a dataset-dependent optimum.
(b) Energy distance versus PCA dimensionality for MixFlow and CFM. MixFlow maintains stable performance while CFM degrades in higher dimensions.
Figure 3: Ablation studies on transcriptomic (Combo-SciPlex) and morphological (BBBC021) datasets. (a) Effect of the number of GMM modes on generation quality. (b) Comparison of MixFlow and CFM robustness across increasing data dimensionality.
7 Limitations and Future Work
While MixFlow demonstrates strong performance across diverse perturbation modalities, several limitations remain. The framework depends on the availability of high-quality perturbation descriptors such as gene embeddings or chemical fingerprints; incomplete or noisy metadata may hinder predictive accuracy. Our evaluation focuses primarily on single perturbations and small combinations, whereas scaling to higher-order combinations, dose–response landscapes, and time-course dynamics remains an open challenge. Moreover, although we provide theoretical justification for out-of-distribution generalisation, the analysis assumes smooth transport fields that may not fully capture heterogeneous tissues or disease states.
Future work should address these challenges by incorporating richer multimodal descriptors, extending the model to capture interaction effects and temporal responses, and scaling training to very large datasets. In particular, applying MixFlow to resources such as the JUMP- consortium dataset for imaging assays (Chandrasekaran et al., 2024) and large single-cell perturbation compendia (Zhang et al., 2025a) could enable the development of true zero-shot models that leverage perturbation similarity at scale. Finally, extending the framework to translational settings such as patient-derived cells or organoids will be critical for realising its potential impact in therapeutic design.
Acknowledgments
This work is supported by Open Targets (Drug2Cell) and by the Chan Zuckerberg Initiative DAF Single-Cell Biology Data Insights program.
Conflict of interest
M.L. has equity interests in Relation Therapeutics, is a scientific co-founder and part-time employee of AIVIVO, and serves on the scientific advisory board of Novo Nordisk.
Code availability
Code is publicly available on github: https://github.com/Lotfollahi-lab/MixFlow.
Appendix A Generalisation analysis, details and proofs
As outlined in Sec. 4, to circumvent the infinite-dimensional Wasserstein manifold of measures and to enable the adoption of theoretical generalization guarantees for finite-dimensional spaces, we assume that the target distribution $\rho$ is itself a Gaussian mixture model (GMM) with $J$ components, fixed mode locations that cover $\mathbb{R}^{D}$ as $J\rightarrow\infty$, and varying mixture weights. Moreover, recall that $\mu$ and $\rho$ denote the base and target distributions, respectively.
Definition 2 (Mixture Wasserstein Distance).
Let $\mu_{GMM(\bm{\Theta},\bm{p})}$ and $\rho_{GMM(\bm{\Gamma},\bm{q})}$ be two measures induced by their corresponding GMMs with $I$ and $J$ modes, mode positions $\bm{\Theta}\in\mathbb{R}^{I\times D}$ and $\bm{\Gamma}\in\mathbb{R}^{J\times D}$, and mode weights $\bm{p}\in\mathbb{S}^{I-1}$ and $\bm{q}\in\mathbb{S}^{J-1}$. The mixture Wasserstein distance between $\mu_{GMM(\bm{\Theta},\bm{p})}$ and $\mu_{GMM(\bm{\Gamma},\bm{q})}$ is defined as follows:
[…]
s.t. $\;\sum_{j=1}^{J}v_{ij}=1\;\;,\;\;\forall i\in\{1,...,I\}$ (9)
$\;\;q_{j}=\sum_{i=1}^{I}v_{ij}\;p_{i}\;\;,\;\;\forall j\in\{1,....,J\}\;\;\;\text{(boundary condition)}$ (10)
$\;\;v_{ij}\geq 0\;\;,\;\;\forall i\in\{1,...,I\},\;\forall j\in\{1,...,J\}$ (11)
This is equivalent to discrete optimal transport between the GMMs’ modes with the transportation cost determined by their mode locations. Intuitively, $v_{ij}$ is the portion of mass of the $i$-th mode in the 1st GMM $\mu$ which is moved to the $j$-th mode in the 2nd GMM $\rho$.
Definition 3 (Mixture Wasserstein Flow).
The matrix $\bm{V}^{I\times J}$ in Eq. 8 induces a time-dependant flow $v:\mathbb{R}^{D}\times[0,1]\rightarrow\mathbb{R}^{D}$} referred to as mixture Wasserstein flow.
Lemma 4 (Dual of Mixture Wasserstein Distance).
The dual form of the constrained optimisation of Eq. 8 is as follows:
$\underset{\bm{z}\in\mathbb{R}^{I+J}}{sup}\;\sum_{i=1}^{I}z_{i}+\sum_{j=1}^{J}q_{j}\;z_{j+I}$ (12)
[…]
$\;\;\{z_{1},z_{2},...,z_{I+J}\}\;\;\;\text{unrestricted in sign , i.e. $\in\mathbb{R}$}$ (14)
Definition 5 (Subset sum condition).
In the discrete optimal transport problem of Eq. 8, the marginal measures $p_{1},...,p_{I}]$ and $q_{1},...,q_{J}]$ are said to meet the subset sum condition if for any proper non-empty subsets of them - denoted by $\{p_{i1},...,p_{i_{I^{\prime}}}\}$ and $\{q_{j1},...,q_{j_{J^{\prime}}}\}$- we have that
$(p_{i1}+...+p_{i_{I^{\prime}}})\neq(q_{j1}+...+q_{j_{J^{\prime}}}).$ (15)
Statement 6.
In the discrete optimal transport problem of Eq. 8 assume that the marginal distributions meet the subset sum condition. Then we have that:
  1. 1.
    Any optimal solution of Eq. 8, $\bm{V}^{*}\in\mathbb{R}^{I\times J}$, has exactly I+J1I+J-1 non-zero elements.
  2. 2.
    A dual constraint in Eq. 13 is binding if and only if the corresponding $\bm{V}^{*}[i,j]=v^{*}_{ij}$ is non-zero.
  3. 3.
    The dual problem of Eq. 12 has a unique optimum $\bm{z}^{*}\in\mathbb{R}^{I+J}$.
The importance of Statement 6 is that - as we will prove in Statement 8 - for any $\mu$ and $\rho$ the subset sum condition, and hence 1, 2, and 3 explained above hold almost surely. Therefore, we can generally assume 1, 2, and 3 in Statement 6 hold for any $\mu$ and $\rho$.
Statement 7.
In the discrete optimal transport problem of Eq. 8 assume that the marginal distributions meet the subset sum condition. Moreover, assume that $\bm{\Theta}$, $\bm{\Gamma}$, $\bm{p}$, and $z^{*}_{1},...,z^{*}_{I}]$ in the dual form are known and fixed. Then $\bm{V}^{*}_{\geq 0}$ is uniquely determined, where $\bm{V}^{*}\in\mathbb{R}^{I\times J}$ is an optimal solution for Eq. 8, and $\bm{V}^{*}_{\geq 0}\in\mathbb{R}^{I\times J}$ is a matrix where the element in its $i$-th row and $j$-th column is 1 if $v^{*}_{ij}>0$ and is 0 otherwise.
Proof: The constraint of Eq. 12 rewrites as
[…]
To maximise the objective of Eq. 12, the variables $z_{j+I}$ should take their maximum possible value. Therefore we have that
[…]
In other words, given $z^{*}_{1},...,z^{*}_{I}]$ the above equation uniquely specifies the values for $z^{*}_{I+1},...,z^{*}_{I+J}]$ and regardless of the probability values of the target distribution $\bm{q}\in\mathbb{R}^{J}$. According to Statement 6 , in Eq. 17 equality holds if and only if $v^{*}_{ij}>0$, and this determines $\bm{V}^{*}_{\geq 0}$. $\blacksquare$
The importance of Statement 7 is that, intuitively, having predicted the closest GMM to $\rho$ (i.e. $\bm{\Theta}$ and $\bm{p}$), having predicted the first $I$ optimal dual values $z^{*}_{1},..,z^{*}_{I}]\in\mathbb{R}^{I}$, and having fixed the GMM modes that cover the feature space (i.e. $\bm{\Gamma}$), $\bm{V}^{*}_{\geq 0}$ is uniquely identified. This is important, because to obtain a predictor $\bm{V}^{*}_{\geq 0}\approx h_{\bm{V}^{*}_{\geq 0}}(\bm{y})$ where $h_{\bm{V}^{*}_{\geq 0}}:\mathbb{R}^{dim(\bm{y})}\rightarrow\{0,1\}^{I\times J}$, one can instead train the following three predictors, and $h_{\bm{V}^{*}_{\geq 0}}(.)$ will be uniquely determined:
  1. 1.
    $\bm{\Theta}\approx h_{\bm{\Theta}}(\bm{y})$, where $h_{\bm{\Theta}}:\mathbb{R}^{dim(\bm{y})}\rightarrow\mathbb{R}^{I\times D}$.
  2. 2.
    $\bm{p}\approx h_{\bm{p}}(\bm{y})$, where $h_{\bm{p}}:\mathbb{R}^{dim(\bm{y})}\rightarrow\mathbb{S}^{I-1}$.
  3. 3.
    $z^{*}_{1},...,z^{*}_{I}]\approx h_{z}(\bm{y})$, where $h_{z}:\mathbb{R}^{dim(\bm{y})}\rightarrow\mathbb{R}^{I}$
Statement 8.
Let $\mu_{GMM(\bm{\Theta},\bm{p})}$ and $\rho_{GMM(\bm{\Gamma},\bm{q})}$ be two measures induced by their corresponding GMMs. Then the marginal distributions $\bm{p}$ and $\bm{q}$ meet the subset sum condition almost surely. Therefore, Statements 6 and 7 hold almost surely.
Proof: There are $(2^{I}-2)\times(2^{J}-2)$ pairs of proper non-empty subsets from the marginal measures $p_{1},...,p_{I}]$ and $q_{1},...,q_{J}]$. Each such a pair like $\{p_{i1},...,p_{i_{I^{\prime}}}\}$ and $\{q_{j1},...,q_{j_{J^{\prime}}}\}$ specify the following subspace of $\mathbb{S}^{I-1}\times\mathbb{S}^{J-1}$ where the subset sum condition is violated:
$p_{i1}+...+p_{i_{I^{\prime}}}=q_{j1}+...+q_{j_{J^{\prime}}}$ (18)
Each such subspace has zero measure, so the union of these $(2^{I}-2)\times(2^{J}-2)$ subspaces has also zero measure. $\blacksquare$
Definition 9 (Projection to subspace of GMMs with $I$ modes).
Let $\rho_{GMM(\bm{\Gamma},\bm{q})}$ be a measure induced by a GMM with modes $\bm{\Gamma}\in\mathbb{R}^{J\times D}$ and mode probabilities $\bm{q}\in\mathbb{S}^{J-1}$. The goal is to find the closest measure to $\rho_{GMM(\bm{\Gamma},\bm{q})}$ which is induced by a GMM with $I$ modes, where $I<<J$ and in terms of mixture Wasserstein distance. This GMM $\mu_{GMM(\bm{\Theta}*,\bm{p}*)}$ is found by:
$\bm{\Theta}^{*},\bm{p}^{*}=\underset{\begin{subarray}{c}\bm{\Theta}\in\mathbb{R}^{I\times D}\\ \bm{p}\in\mathbb{R}^{I}\end{subarray}}{argmin}\;\;\mathcal{M}\mathcal{W}^{2}_{2}(\mu_{GMM(\bm{\Theta},\bm{p})},\rho_{GMM(\bm{\Gamma},\bm{q})})$ (19)
Let’s assume we have the 3 predictors $h_{\bm{\Theta}}(.)$, $h_{\bm{p}}(.)$, and $h_{z}(.)$ explained above, that take in the population descriptor $\bm{y}$ and predict the projection of the unseen $\rho$ to the space of GMMs (i.e. $\bm{\Theta}^{*}$ and $\bm{p}^{*}$) as well the $I$ dual solution values $z^{*}_{1},...,z^{*}_{I}]$. Given these, according to Statement 7 we can uniquely determine I+J1I+J-1 elements of $\bm{V}^{*}\in\mathbb{R}^{I\times J}$ which are non-zero, and the rest are uniquely set to zero. Recall that the goal is to predict the flow $\bm{V}^{*}\in\mathbb{R}^{I\times J}$, because when it is applied to the measure $\mu_{GMM(\bm{\Theta}*,\bm{p}*)}$ it gives us the prediction for $\rho$. Having identified non-zero elements of $\bm{V}^{*}$, Statements 10 and 11 impose linear constraints on the identified elements, thereby improving the idenfication and prediction of $\bm{V}^{*}$.
Statement 10 (Barycentre condition).
For the matrix $\bm{\Theta}^{*}\in\mathbb{R}^{I\times D}$ in Eq. 19 we have that
$\bm{\Theta}^{*}_{i}=\sum_{j=1}^{J}v^{*}_{ij}\bm{\Gamma}_{j}.$ (20)
Proof: We compute the derivative of Eq. 8 with respect to $\bm{\Theta}_{i}$ and set it to zero:
[…]
$\frac{\partial}{\partial\bm{\Theta}_{i}}\sum_{j=1}^{J}\big[\bm{\Theta}_{i}^{T}\bm{\Theta}_{i}-2\bm{\Theta}_{i}^{T}\bm{\Gamma}_{j}\big]\;v^{*}_{ij}\;p_{i}=$ (22)
$\sum_{j=1}^{J}[2\bm{\Theta}_{i}-2\bm{\Gamma}_{j}]v^{*}_{ij}p_{i}=$ (23)
$2p_{i}\;[\bm{\Theta}_{i}\cancelto{1}{\sum_{j=1}^{J}v^{*}_{ij}}-\sum_{j=1}^{J}v^{*}_{ij}\bm{\Gamma}_{j}]=0$ (24)
$\Rightarrow$ $\bm{\Theta}_{i}^{*}=\sum_{j=1}^{J}v^{*}_{ij}\bm{\Gamma}_{j}$ (25)
$\blacksquare$
Statement 11 (Optimality condition for mode probabilities).
The following optimality condition
$\frac{\partial}{\partial\bm{p}}\mathcal{M}\mathcal{W}^{2}_{2}(\mu_{GMM(\bm{\Theta}^{*},\bm{p})},\rho_{GMM(\bm{\Gamma},\bm{q})})|_{\bm{p}=\bm{p}^{*}}=0$ (26)
implies that
[…]
Proof: At the optimal point $\bm{\Theta}^{*},\bm{p}^{*}]$, if two mode probabilities like $p^{*}_{i_{1}}$ and $p^{*}_{i_{2}}$ are changed by $\epsilon$ and $\epsilon$ (where $\epsilon$ is a small positive scalar), we have that
$\frac{change\;in\;\mathcal{M}\mathcal{W}^{2}_{2}}{change\;in\;\bm{p}}$ (28)
$=\frac{1}{\epsilon}\Big[\sum_{j=1}^{J}||\bm{\Theta}^{*}_{i_{1}}$- $\bm{\Gamma}_{j}||_{2}^{2}\;v^{*}_{i_{1}j}\;(+\epsilon)\Big]+\frac{1}{\epsilon}\Big[\sum_{j=1}^{J}||\bm{\Theta}^{*}_{i_{2}}-\bm{\Gamma}_{j}||_{2}^{2}\;v^{*}_{i_{2}j}\;(-\epsilon)\Big]$ (29)
$=\Big[\sum_{j=1}^{J}||\bm{\Theta}^{*}_{i_{1}}$- $\bm{\Gamma}_{j}||_{2}^{2}\;v^{*}_{i_{1}j}\Big]-\Big[\sum_{j=1}^{J}||\bm{\Theta}^{*}_{i_{2}}-\bm{\Gamma}_{j}||_{2}^{2}\;v^{*}_{i_{2}j}\Big]=0$ (30)
Therefore
[…]
$\blacksquare$
Now we are ready to state the training phase and testing phase of the learning algorithm.
Statement 12 (Training phase).
Given is a set of $N$ GMM-induced measures $\{\rho_{GMM(\bm{\Gamma},\bm{q}^{(n)})}\}_{n=1}^{N}$ and their corresponding population descriptors $\{\bm{y}_{1},..,\bm{y}_{N}\}$. Note that the matrix $\bm{\Gamma}\in\mathbb{R}^{J\times D}$ are the GMM modes in the target distribution that carpet the feature space, and are kept fixed during trianing and testing, hence to superscript $n$ for $\bm{\Gamma}$. For each dataset instance, the following vectors and matrices are obtained
  1. 1.
    The parameters of the closest GMM to $\rho_{GMM(\bm{\Gamma},\bm{q}^{(n)})}$ and with $I$ modes:
    $\bm{\Theta}^{*(n)},\bm{p}^{*(n)}=\underset{\begin{subarray}{c}\bm{\Theta}\in\mathbb{R}^{I\times D}\\ \bm{p}\in\mathbb{R}^{I}\end{subarray}}{argmin}\;\;\mathcal{M}\mathcal{W}^{2}_{2}(\mu_{GMM(\bm{\Theta}^{,}\bm{p})},\rho_{GMM(\bm{\Gamma},\bm{q}^{(n)})})$ (32)
  2. 2.
    Having obtained $\bm{\Theta}^{*(n)}$ and $\bm{p}^{*(n)}$, solve the dual form of Eq. 12 for the distributons $\mu_{GMM(\bm{\Theta}^{*(n)},\bm{p}^{*(n)})}$ and $\rho_{GMM(\bm{\Gamma}^{(n)},\bm{q}^{(n)})}$. Save the first $I$ elements of the dual solution $z^{*(n)}_{1},...,z^{*(n)}_{I}$].
Now having obtained $\{\bm{\Theta}^{*(n)},\bm{p}^{*(n)},[z^{*(n)}_{1},...,z^{*(n)}_{I}]\}_{n=1}^{N}$, the following predictors are trained
  1. 1.
    $\bm{\Theta}^{*(n)}\approx h_{\bm{\Theta}}(\bm{y}_{n})$, where $h_{\bm{\Theta}}:\mathbb{R}^{dim(\bm{y})}\rightarrow\mathbb{R}^{I\times D}$.
  2. 2.
    $\bm{p}^{*(n)}\approx h_{\bm{p}}(\bm{y}_{n})$, where $h_{\bm{p}}:\mathbb{R}^{dim(\bm{y})}\rightarrow\mathbb{S}^{I-1}$.
  3. 3.
    $z^{*(n)}_{1},...,z^{*(n)}_{I}]\approx h_{z}(\bm{y}_{n})$, where $h_{z}:\mathbb{R}^{dim(\bm{y})}\rightarrow\mathbb{R}^{I}$
Statement 13 (Testing phase).
In the testing phase, a population descriptor $\bm{y}_{test}$ is given. Firstly, the following vectors/matrices are predicted:
  1. 1.
    $\bm{\Theta}^{(test)}\approx h_{\bm{\Theta}}(\bm{y}_{test})$.
  2. 2.
    $\bm{p}^{(test)}\approx h_{\bm{p}}(\bm{y}_{test})$
  3. 3.
    $z^{(test)}_{1},...,z^{(test)}_{I}]\approx h_{z}(\bm{y}_{test})$
Now having obtained $\big[\bm{\Theta}^{(test)},\bm{p}^{(test)},[z^{(test)}_{1},...,z^{(test)}_{I}$$\big]$, according to Statement 7 $\bm{V}^{test}_{\geq 0}$ is uniquely identified and according to Statement 8 almost surely has exactly I+J1I+J-1 non-zero elements. In other words, at this point we are left with I+J1I+J-1 variables (i.e. $v^{test}_{ij}$-s) that we want to identify as closely as possible. For these I+J1I+J-1 variables, we have the following constraints
  1. 1.
    Each barycentre condition in Eq. 20 provides $D$ linear constraints on non-zero $v^{test}_{ij}$-s, resulting in $I\times D$ linear constraints. Note that the matrix $\bm{\Gamma}\in\mathbb{R}^{J\times D}$ are the GMM modes in the target distribution that carpet the feature space, and are kept fixed throughout.
  2. 2.
    Each constraint of Eq. 27 provides a linear constraint, resulting in II more linear constraints. We add the constant scalar in Eq. 27 to our variable list, so we have $I+J$ variables to identify.
In sum, we have $I+J$ variables, $ID+I$ constraints, and hence JIDJ-ID degrees of freedom. Having estimated $\bm{\Theta}^{(test)}$ and the flow $\bm{V}^{(test)}$, we can apply the latter on the former to arrive at an estimate for $\rho^{(test)}$
Statement 14.
If $I=1$ the dual problem of Eq. 12 is ill-defined. Proof: If $I=1$, according to Eq. 17, we have that
$z^{*}_{j+I}=\||\bm{\Theta}_{1}-\bm{\Gamma}_{j}||_{2}^{2}-\frac{z^{*}_{1}}{\cancelto{1}{p_{1}}}.$ (33)
If we substitute this in Eq. 12
$\underset{\bm{z}\in\mathbb{R}^{1+J}}{sup}\;$ $z_{1}+\sum_{j=1}^{J}q_{j}\;z_{j+I}=$ (34)
$\underset{\bm{z}\in\mathbb{R}^{1+J}}{sup}\;$ $z_{1}+\sum_{j=1}^{J}q_{j}\;\Big[\||\bm{\Theta}_{1}-\bm{\Gamma}_{j}||_{2}^{2}-\frac{z_{1}}{\cancelto{1}{p_{1}}}\Big]=$ (35)
$\underset{\bm{z}\in\mathbb{R}^{1+J}}{sup}\;$ $z_{1}-z_{1}\cancelto{1}{\sum_{j=1}^{J}q_{j}}+\sum_{j=1}^{J}q_{j}\;\Big[\||\bm{\Theta}_{1}-\bm{\Gamma}_{j}||_{2}^{2}\Big]$ (36)
In the last equation $z_{1}$ is cancelled out from the objective. $\blacksquare$
Essentially, according to Statement 14 the described training and testing procedure fail, and this discourages the use of uni-modal Gaussian distribution as the base distribution.